Let the six-month interest rate be 3% in UK, the spot rate be $1.60/₤, and the six-month forward rate be $1.61/₤. Assume the IRP condition holds. Find the six-month interest rate in the U.S.
Answer #1

6-month Interest Rate in UK = 3%

Spot Rate = $1.60/₤

6-month Forward rate = $1.61/₤

\frac{Forward Rate}{Spot rate} = (\frac{1+US Interest Rate/2}{1+ UK InterestRate/2 })

\frac{1.61}{1.60} = (\frac{1+US Interest Rate/2}{1+ (0.03/2)})

1+ US Interest rate/2= 1.02134375

US Interest rate = 4.26875%

So, 6-month US interest rate is 4.27%